OpenStudy (anonymous):

square root of -147 + the square root of 192 - the square root of 243

OpenStudy (anonymous):

\[\sqrt{-147}+\sqrt{192}-\sqrt{243}=\sqrt{7*7*-3}+\sqrt{8^{2}*3}-\sqrt{3*3*3*3*3}\] \[=7\sqrt{3}\iota +8\sqrt{3}-9\sqrt{3}\] \[=7\sqrt{3}\iota-\sqrt{3}\]